OpenStudy (anonymous):

how many atoms of oxygen are in 0.100 mol of silicon dioxide? please explain how to get answer!

OpenStudy (anonymous):

SiO2 n=N/L -> N= n*L = 0,1 * 6,022*10^23= 6,022*10^22 molecules of SiO2 *2 equals atoms of O = 1,2044*10^23

OpenStudy (anonymous):

There are 2 atoms of oxygen in \[SiO _{2}\] . Therefore, using Avogadro's number (which tells you the number of particles per mol of any given substance) and dimensional analysis: 0.100 mol SiO_2 | 6.022 * 10^23 | 2 atoms | ------------------------------------------ = 1.20 * 10^23 atoms | mol SiO_2 |

OpenStudy (anonymous):

Thanks so much!